OpenStudy (anonymous):

Which of the following describes the transformation of the graph y = x^2 in graphing y = -x^2 - 5? reflect over the x-axis and shift down 5 reflect over the y-axis and shift down 5 reflect over the x-axis and shift left 5

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

last one?

OpenStudy (perl):

$$ \Large{ y = x^2 \\ \to y=-x^2 ~~~~\text{ flip about x axis} \\ \to y =-x^2 - 5 ~~~\text{ shift down 5 units} }$$

OpenStudy (anonymous):

thanks